OpenStudy (anonymous):

If f(x) = x^2 + 1 and g(x) = 3x + 1, find [f(2)-g(1)]^2

OpenStudy (anonymous):

When you post a question and go off line, you are indicating that you just want the answer and are not here to learn. So I was forced to delete my posted help. Holler when you need help solving this and I will be glad to assist.

OpenStudy (anonymous):

That's very condusive to learning ehuman, to think someone might have other things going on throughout the day that may impede them from staying logged in is ludicrous.

OpenStudy (lucaz):

\[f(2)=subtitute~x~in~f(x)~by~2=(2)^2\] \[same~process~for~g(1)=3(1)+1\]

OpenStudy (loser66):

to support lucaz, f(2) = substitute x inf(x)by 2 = (2)^2 +1

OpenStudy (anonymous):

So the answer is 5?

OpenStudy (loser66):